# Compact Speech Translation Models via Discrete Speech Units Pretraining

### v1

@@ Abstract
Using Self-Supervised Learning (SSL) as model initialization is now common to obtain strong results in Speech Translation (ST). However, they also impose a large memory footprint, hindering on-device deployment. In this paper, we leverage the SSL models by pretraining smaller models on their Discrete Speech Units (DSU). We pretrain encoder-decoder models on 1) Filterbank-to-DSU and 2) DSU-to-Translation data, and take the encoder from 1) and the decoder from 2) to initialise a new model, finetuning this on limited speech-translation data. The final model becomes compact by using the DSU pretraining to distil the knowledge of the SSL model. Our method has several benefits over using DSU as model inputs, such as shorter inference pipeline and robustness over (DSU) tokenization. In contrast to ASR pretraining, it does not require transcripts, making it applicable to low-resource settings. Evaluation on CoVoST-2 X-En shows that our method is >$0.5$ BLEU better than a ST model that directly finetune the SSL model, given only half the model size, and on a par with ASR pretraining.
@@ 1 Introduction
In Speech-to-Text Translation (ST), the use of SSL models, e.g., wav2vec2 and HuBERT Baevski et al. (2020); Hsu et al. (2021), as model initialization is now common to obtain SOTA result Agarwal et al. (2023). Nevertheless, using these bulky models as initialisation imposes a large memory footprint to the downstream ST, which limits flexibility in modelling choices. In addition, it hinders on-device deployment that is crucial for privacy and useful in the absence of internet connection.
Knowledge distillation (KD) Kim and Rush (2016); Inaguma et al. (2021) is the standard approach to creating smaller models, but obtaining the high-quality pseudo labels for training ST systems is very costly. Instead, we focus on Discrete Speech Units (DSU) since they can be obtained cheaper, e.g., without requiring a translation model, and are an intermediate representation between speech and text. DSUs are K-means clusters of speech representations from selected layers within a SSL model. It represents sequences of discrete tokens, which are easier to model within a text processing architecture Polyak et al. (2021); Chou et al. (2023). DSU sequences are far smaller than dense Filterbank sequences, and they can be shortened with removing duplicates and Byte Pair Encoding (BPE) Sennrich et al. (2016). Therefore, in ST, a straightforward method to distill the SSL models is to replace the dense Filterbank by the DSUs, aka the DSU-to-Translation (DSU-to-Trl) model. Although using DSUs for training allows for transfer learning, using them at inference requires large and slow models.
To address the above, we use DSU for pretraining rather than as model input to make ST models more compact. Our method leverages large SSL models by pretraining smaller models on the corresponding DSU. More specifically, it pretrains encoder-decoder models on 1) Filterbank-to-DSU (Fbk-to-DSU) and 2) DSU-to-Trl data, and takes the encoder from 1) and the decoder from 2) to initialise a new model. After that, the new model is finetuned on limited speech-translation data. Under this formulation, (1) the model is much more compact than the SSL model (2) the compact model does not use DSU as inputs which (a) avoids the lengthy SSL and clustering inference pipeline and (b) minimizes its sensitivity to (DSU) tokenization. (3) Our method does not require transcripts, unlike ASR pretraining, making it applicable to low-resource settings. Similar to ST methods that use pretrained components, our method could be limited by the pretraining modality gap Liu et al. (2020); Le et al. (2023). Motivated by prior works, we investigate mitigating it with Connectionist Temporal Classification (CTC) Graves et al. (2006) regularisation.
We evaluate our method on CoVoST-2 Wang et al. (2021) X-En (21 language directions) with multilingual ST. By simply using a monolingual HuBERT model to extract the DSU, our method has a gain of >$3$ BLEU with respect to training from scratch, demonstrating DSU’s cross-lingual ability. Our method is also >$0.5$ BLEU better than a ST model that directly finetuned HuBERT despite having half model size. Our main contributions are:
  • We propose to distill knowledge from large SSL models to make ST models more compact via DSU pretraining. Our method avoids the lengthy inference pipeline in the DSU-to-Trl method.
  • We examine the effect of (DSU) tokenization on the ST models. Our method is robust across different tokenizations, which is not the case for the DSU-to-Trl method.
  • We examine applying CTC to the DSU in pretraining and on the translation in finetuning to mitigate the modality gap, and we show that both improve performance.
@@ 2 Related Work
There are a number of related works that use DSU to enhance ST. Fang and Feng (2023) and Zhang et al. (2023b) use DSU to create more training data in a back-translation fashion. Chang et al. (2023) and Zhang et al. (2023b) explore the replacement of Filterbank by DSU as speech input. In contrast, we use DSU and its translation model for pretraining, resulting in a better Fbk-to-Trl model that has a shorter inference pipeline.
@@
AST model (#Params) High Mid Low All
Scratch (52M) $19.37$ $7.91$ $0.73$ $5.99$
ASR-Pretraining (52M) $26.53$ $12.22$ $1.82$ $9.00$
Hu-Transformer (113M) $24.31$ $11.42$ $2.18$ $8.60$
DSU-Adapter (48M) $26.54$ $12.91$ $1.76$ $9.13$
Table 1: Results in BLEU on CoVoST-2 X-En. DSU-Adapter is 0.53 points better than Hu-Transformer that is $2.3$ times larger and on a par with ASR-Pretraining.
@@ 3 Method
Our method distills knowledge from the SSL representations via first DSU pretraining (PT) and then finetuning (FT) on the paired ST data.
In PT, our method uses DSU from an SSL speech model to train a pair of encoder-decoder models: 1) Fbk-to-DSU and 2) DSU-to-Trl. The Fbk-to-DSU model takes the log Mel Filterbank as the encoder input and predicts the corresponding DSU sequence. The model is trained by Maximum Likelihood Estimation (MLE). In addition, we interpolate between MLE and CTC applied to the last encoder layer:
$\mathcal{L}^{\text{Fbk-to-DSU}}=(1-\lambda_{\alpha})\mathcal{L}_{\text{MLE}}(\mathbf{U}|\mathbf{F})+\lambda_{\alpha}\mathcal{L}_{\text{CTC}}(\mathbf{\tilde{U}}|\mathbf{F})$ (1)
@@
In FT, we use the encoder of the Fbk-to-DSU model and the decoder (and its output layer) of the DSU-to-Trl model to initialise the compact model, followed by finetuning on the paired speech-translation data using both MLE and CTC loss Gaido et al. (2021); Zhang et al. (2023a):
$\mathcal{L}^{\text{FT}}=(1-\lambda_{\beta})\mathcal{L}_{\text{MLE}}(\mathbf{Y}|\mathbf{F})+\lambda_{\beta}\mathcal{L}_{\text{CTC}}(\mathbf{\tilde{Y}}|\mathbf{F})$ (2)
@@
Tokenization
As mentioned in section 1, we examine the tokenization effect, including using separate (source/target) vocabularies or not. We denote each DSU token by a hashtag with an index, e.g,. ’#456’. Following Ao et al. (2022), we reduce the sequence length via merging the sequential repetitions, e.g, ’#1 #1 #1 #456 #456 #23’ becomes ’#1 #456 #23’. When BPE is applied on the DSU, the whitespace characters are ignored, e.g., ’#1 #456 #23’ becomes ’#1#456#23’ which could be split into ’#1#45 6#2 3’.
@@
Table 2: Tokenization effect on 4 types of ST methods. In general, darker (brighter) cells refer to weaker (stronger) models. The best two models apply both BPE on the DSU and separate vocabularies in PT (cells in yellow).
4 Experiments
4.1 Data Preprocessing
We follow standard practices to preprocess the CoVoST-2 X-En data, i.e., log-Mel Filterbank of $80$ dimensions for speech inputs and BPE with $8$K vocabulary on the English target, $16$K on the transcripts and $32$K on the DSU, unless otherwise specified. We use an English HuBERT-Base model to extract the DSU by first downsampling the CoVoST-2 audio to $16$KHz. Each audio data utterance is then converted into the DSU, i.e., the clustering indexes, by applying K-Means clustering (K=$1,000$) on its HuBERT representation from the 6th layer Lakhotia et al. (2021). We sample training data from each language pair to form a subset of $98$K instances to train the K-Means model. Further details are in Appendix A.1.
4.2 Model Configuration
All models are based on Transformer Vaswani et al. (2017) with implementations from fairseq Ott et al. (2019); Wang et al. (2020). In the Fbk-to-token [^1] Token refers to either transcriptions, DSU, or translations. models, the encoder has convolutional layers to downsample the Fbk by a factor of 4. There are 12-6 layers in the transformer encoder-decoder, whereas the embedding and feed-forward network (FFN) dimensions are 256 and 4,096 respectively, unless otherwise specified. It is worth noting that: (1) The Fbk-to-DSU model is not trained on the translations, so it is not directly comparable to the ST models. Its effect to ST lies on its pretrained encoder (Table 2). (2) The DSU-to-Trl model is a ST model, and it can also be used for initialization.
DSU-to-Trl follows the Transformer used in text translation. We use 6-6 layers in the encoder-decoder which the dimension of embedding and FFN is 256 and 2,048 respectively. Despite its smaller model size, its inference requires storing and calling both HuBERT and K-Means models.
DSU-Adapter is our proposed method. To better align the two pre-trained components, we also experiment with adding an extra encoder layer [^2] All parameters are trainable in ST finetuning. after the pre-trained encoder. Since its decoder is initialized by the DSU-to-Trl method, its decoder FFN dimension is $2,048$.
4.3 Training and Inference
We leave the details in Appendix A.3. It is worth noting that we do not use extra audio data, e.g., Libri-Light Kahn et al. (2020) in pretraining.
5 Results and Analysis
We divide the 21 language pairs by resource level into: 1) "High": {ca, de, es, fr}, 2) "Mid": {fa, it, pt, ru and zh}, 3) "Low": {ar, cy, et, id, ja, lv, mn, nl, sl, sv, ta, tr} and 4)"All": the 21 languages pairs. We report the average BLEU over the test sets of each group using sacrebleu Post (2018)333nrefs:1—case:mixed—eff:no—tok:13a—smooth:exp—version:2.3.1.
Table 1 compares our DSU-Adapter and the baselines. Our DSU-Adapter is $3$ BLEU points (in the group "All") higher than Scratch. Furthermore, it is better than Hu-Transformer in spite of having half the parameters. For "Mid" and "High", the improvement is $1.49$ and $2.23$ points respectively, but it falls short by $0.42$ points for "Low". We also compare to ASR pretraining, which is not always applicable, e.g., in low-resource setting or perhaps even in an unwritten language Zhang et al. (2022a). Surprisingly, our adapter is on a par with it, i.e., $0.13$ points better. The result remains consistent when measured in chrF Popović (2015) and COMET [^4] WMT22-COMET-DA Rei et al. (2022), see Appendix A.4.
5.1 Tokenization and its Transfer
Table 2 presents the effect of tokenization. The 1st column indicates if BPE is applied on the DSU, and the 2nd column "$|\mathcal{V}|$" shows the vocabulary size, and if the model has a joint vocabulary or not in PT. For example, "$1$K-$8$K" means that we use a vocabulary of size $1$K for DSU and $8$K for English. These results are for all languages, please see Appendix A.4 for fine grained results.
Effect of BPE on DSU-to-Trl

### v2

@@ Abstract
We propose a pretraining method to use Self-Supervised Speech (SSS) model to creating more compact Speech-to-text Translation. In contrast to using the SSS model for initialization, our method is more suitable to memory constrained scenario such as on-device deployment. Our method is based on Discrete Speech Units (DSU) extracted from the SSS model. In the first step, our method pretrains two smaller encoder-decoder models on 1) Filterbank-to-DSU (Fbk-to-DSU) and 2) DSU-to-Translation (DSU-to-Trl) data respectively. The DSU thus become the distillation inputs of the smaller models. Subsequently, the encoder from the Fbk-to-DSU model and the decoder from the DSU-to-Trl model are taken to initialise the compact model. Finally, the compact model is finetuned on the paired Fbk-Trl data. In addition to being compact, our method requires no transcripts, making it applicable to low-resource settings. It also avoids speech discretization in inference and is more robust to the DSU tokenization. Evaluation on CoVoST-2 (X-En) shows that our method has consistent improvement over the baseline in three metrics while being compact i.e., only half the SSS model size.
@@ 1 Introduction
In Speech-to-text Translation (ST), using Self-Supervised Speech (SSS) models, such as wav2vec 2.0 and HuBERT Baevski et al. (2020); Hsu et al. (2021), as model initialization is now common to obtain the SOTA result Agarwal et al. (2023). Nevertheless, such model initialisation makes the ST model less memory-adaptive and could impose a large memory footprint. These factors hinders on-device deployment that is crucial for privacy and useful in the absence of internet connection.
How can we use the SSS model(s) to create a more compact ST model? When using the SSS model for initialization, the corresponding ST model uses the dense representations of the SSS model for its task. Alternatively, an informative proxy, which requires less memory to obtain, for the dense representation may make the ST model more compact.
Discrete Speech Units (DSU) extracted from the SSS model can be such a good proxy. DSU are K-Means clusters of speech representations from selected layers of the SSS model. It represents sequence of discrete tokens, which are easier to model within a text processing architecture Polyak et al. (2021); Chou et al. (2023). DSU sequences [^1] In this paper, DSU and DSU sequences are used interchangeably. When we need to focus on a few units of the sequence, we call them DSU tokens. are far smaller than the sequences of dense representations. Therefore, a straightforward method to distill the SSS models is to use DSU as speech inputs, aka the DSU-to-Translation (DSU-to-Trl) model. Although using DSU as inputs allows for transfer learning and a memory-adaptive model, using them at inference still requires storing and calling the quantization modules, i.e, the SSS model and the K-Means model.
We thus propose to use DSU for pretraining (PT) rather than as model input to make ST models more compact. Our method distils the SSS model by pretraining smaller models on the corresponding DSU. More specifically, our method firstly pretrains two smaller encoder-decoder models on 1) Filterbank-to-DSU (Fbk-to-DSU) and 2) DSU-to-Trl data respectively. The DSU thus become the distillation inputs of the smaller models. Subsequently, the encoder from the Fbk-to-DSU model and the decoder from the DSU-to-Trl model are taken to initialise the compact model. Finally, the compact model is finetuned on the paired Fbk-Trl data. Under this formulation, (1) we can use the SSS model to create a ST model that is adaptive to the memory footprint. (2) Our method requires no transcripts, unlike ASR-pretraining, making it applicable to low-resource settings. (3) Our method avoids using the quantization modules in inference. (4) Extensive results also show that our method is more robust to DSU tokenization than the DSU-to-Trl method.
We evaluate our method on CoVoST-2 Wang et al. (2021) X-En language directions ($21$ in total) using multilingual ST. By using a HuBERT-Base model to extract the DSU, our method shows strong and consistent improvements in three evaluation metrics with respect to a ST model that is trained from scratch. Our main contributions are:
  • We propose a pretraining method to distil the SSS model to creating a more compact ST model. Rather than competing with the SOTA ST models, adaptability to the memory footprint is our key focus.
  • Our method uses DSU for pretraining rather than as model inputs. This lowers the inference cost, especially for on-device purpose, by avoiding the quantization modules (storage and running).
  • We conduct extensive analysis to study the effect of DSU tokenization to both using DSU as model inputs and as pretraining. Our pretraining method is found to be more robust to different tokenizations.
@@ 2 Related Work
There are a number of related works that use DSU to enhance ST. Fang and Feng (2023) and Zhang et al. (2023b) use DSU to create more training data in a back-translation fashion. Chang et al. (2023) and Zhang et al. (2023b) explore the replacement of Filterbank by DSU as speech input. Furthermore, Yan et al. (2024) proposes a multi-tasking learning framework with hard parameter sharing, i.e., using a joint vocabulary for text tokens and DSU, to improve the speech-text modality gap. In contrast, we use DSU and its translation model for pretraining, resulting in a better Fbk-to-Trl model that has a shorter inference pipeline.
@@ 3 Method
Figure 1: Illustration of the Fbk-to-DSU model. It is like an auto-encoding training process, but between a continuous format (log Mel Filterbank) and its discrete format (DSU) that is extracted from a HuBERT model.
Our method uses DSU in the form of pretraining to distil knowledge from the SSS (dense) representations to creating a more compact ST model.
In the first step, our method pretrains two smaller encoder-decoder models on 1) Fbk-to-DSU and 2) DSU-to-Trl data respectively. The Fbk-to-DSU model takes the log Mel Fbk as the encoder input and predicts the DSU sequence. The model is trained by an interpolation of Connectionist Temporal Classification (CTC, Graves et al. (2006)) loss that is applied to the last encoder layer and label-smoothed Cross-Entropy (CE) loss:
$\mathcal{L}^{\text{Fbk-to-DSU}}=(1-\lambda_{\alpha})\mathcal{L}_{\text{CE}}(\mathbf{U}|\mathbf{F})+\lambda_{\alpha}\mathcal{L}_{\text{CTC}}(\mathbf{\tilde{U}}|\mathbf{F})$ (1)
@@
Next, we use the encoder of the Fbk-to-DSU model and the decoder (and its output layer) of the DSU-to-Trl model to initialise the compact model, followed by finetuning on the paired Fbk-Trl data using both CE and CTC loss Gaido et al. (2021); Zhang et al. (2023a) on the translations:
$\mathcal{L}^{\text{FT}}=(1-\lambda_{\beta})\mathcal{L}_{\text{CE}}(\mathbf{Y}|\mathbf{F})+\lambda_{\beta}\mathcal{L}_{\text{CTC}}(\mathbf{\tilde{Y}}|\mathbf{F})$ (2)
@@
3.1 Tokenization of DSU in different models
Figure 2: Aligning the DSU tokenization of the Fbk-to-DSU, DSU-to-Trl and compact ST model.
The discrete nature of DSU makes the above training process similar to the transcripts-based pretraining. However, DSU is self-supervised, whereas transcripts require human annotations. DSU are also much longer and can be represented with various sets of symbols.
The length issue could be relieved by merging sequential repetitions Ao et al. (2022), e.g., ’#$1$$1$$1$$456$$456$$23$’ becomes ’#$1$$456$$23$’, where each DSU token is denoted by a #{integer}. Byte Pair Encoding (BPE) Sennrich et al. (2016) could be applied to reduce the DSU sequence length further, e.g., ’#$1$$456$$23$’ could be split into a single subword unit: ’#$1$#$456$#$23$’.
Since both Fbk-to-DSU and DSU-to-Trl models map to different targets, and DSU can be represented with various set of symbols, we align the tokenizations (or called vocabularies [^2] We use vocabulary and tokenization interchangeably, since we did not apply subword regularisation.) of the two models. Figure 2 provides an illustration. The vocabulary of the Fbk-to-DSU model (Vocabulary A) is identical to the source vocabulary of the DSU-to-Trl model (their weights are not shared since these two models are trained independently), whereas the target vocabulary (Vocabulary B) of the DSU-to-Trl model is identical to the target vocabulary of the final compact model (their weights are shared during initialisation). The DSU-to-Trl model is similar to a text translation model, so we also experiment of using separate vocabularies or a joint vocabulary. If a joint vocabulary of English subword units and DSU (BPE or not) is used, all the three models would have the same vocabulary, and the weights of the source and target vocabularies of the DSU-to-Trl model are also tied.
4 Experiments
4.1 Data Preprocessing
We follow standard practices to preprocess the CoVoST-2 X-En data. For speech inputs using $80$-D log Mel Fbk, we computed the features for every $10$ms with a $25$ms window and then normalized them using its mean and variance computed over each channel. We use the BPE implementation from sentencepiece Kudo and Richardson (2018) and obtain vocabulary of size $8$K on the English target, $16$K on the (non-English) transcripts and $32$K on the DSU, unless otherwise specified.
We use HuBERT-Base [^3] https://github.com/facebookresearch/fairseq/tree/main/examples/textless_nlp/gslm/speech2unit model to extract the DSU by first downsampling the CoVoST-2 audio to $16$KHz. Each audio data utterance is then converted into the DSU, i.e., the clustering indexes, by applying K-Means clustering (K=$1,000$; MiniBatchKMeans from sklearn) on its HuBERT representation from the $6$th layer Lakhotia et al. (2021). To train the K-Means model, we divide the $21$ language pairs into three groups: 1) {ar, cy, et, id, ja, lv, mn, sl, sv, ta, tr}, 2) {nl, pt, ru, zh} and 3) {ca, de, es, fa, fr, it}. We then sample $1$K instances for each language pair in group 1), which becomes $3$K in group 2) and $12.5$K in group 3), to create a multilingual training dataset of $98$K instances for the K-Means model.
4.1.1 On the choice of using HuBERT-Base
Given the rapid advance in the SSS models, there are many alternatives, such as XLS-R Babu et al. (2021) and Wavlm Chen et al. (2022), for extracting the DSU for our method. These models are larger in scale and could be multilingual, thus providing DSU of higher qualities. The improvement of our method by using DSU from the HuBERT-Base would probably be a lower-bound, considering its relatively poor qualities to the bigger models. Since our goal is about compactness via DSU pretraining rather than comparing the DSU qualities across the SSS models, we took a simple HuBERT-Base model to illustrate the idea. Pretraining only on English audio data could also suggest hints on whether the DSU and our method could be generalised to languages that are unseen to the SSS models.
4.2 Model Configuration
All models are based on Transformer Vaswani et al. (2017) with implementations from fairseq Ott et al. (2019); Wang et al. (2020). In the Fbk-to-Token (i.e. transcriptions, DSU, or translations) models, the encoder has convolutional layers to downsample the Fbk by a factor of $4$. There are $12$-$6$ layers in the transformer encoder-decoder, whereas the embedding and feed-forward network (FFN) dimensions are $256$ and $4,096$ respectively, unless otherwise specified. It is worth noting that: (1) The Fbk-to-DSU model is not trained on the translations, so it is not directly comparable to the ST models. Its effect on ST lies on its pretrained encoder (Table 2). (2) The DSU-to-Trl model is a ST model which decoder can be used for initialization.
DSU-to-Trl follows the Transformer used in text translation. We use $6$-$6$ layers in the encoder-decoder which the dimension of embedding and FFN is $256$ and $2,048$ respectively. In addition, we use "pre" layer-normalization Nguyen and Salazar (2019). Despite its smaller model size, its inference requires the quantization modules.
DSU-Adapter is our proposed method. To better align the two pre-trained components, we also experiment with adding an extra encoder layer as a simple adapter layer after the pre-trained encoder. Because of the small model size, all model parameters are trainable. Since its decoder is initialized by the DSU-to-Trl method, its decoder FFN dimension is $2,048$.
4.3 Training and Inference
It is worth noting that we do not use extra audio data, e.g., Libri-Light Kahn et al. (2020) in our (pretraining) experiments. Furthermore, we apply the following conditions in (pre-)training:
  • We skip training data that are longer than $30$ seconds (audio) or $1,024$ target tokens.
  • We set $\lambda_{\alpha}$ and $\lambda_{\beta}$ in CTC to $0.3$ and the smoothing parameter to $0.1$
  • We use Adam optimizer with inverse square root scheduler for all model training.
  • In all Fbk-to-Token models, the effective mini-batch size, warm-up steps, peak learning rate and training steps are $32$K frames, $25$K, $2$e33-3- 3 and $60$K steps respectively.
  • Similarly, in all DSU-to-Trl models, we use $80$K tokens, $10$K, $5$e44-4- 4 and $50$K steps.
  • Similarly, in Hu-Transformer, we use $4$M frames, $4$K, $1$e44-4- 4 and $300$K steps.
In inference, we average the last 5 checkpoints and use beam size of $5$ in generation. All experiments are run on Nvidia A100 GPUs. It takes about $1$ day for $2$ A100 ($40$GB) GPUs to complete an experiment that uses Filterbank as speech inputs.
AST model (#Params) BLEU chrF COMET-22-DA
High Mid Low All High Mid Low All High Mid Low All
Scratch (52M) $19.4$ $7.91$ $0.73$ $5.99$ $43.6$ $27.2$ $14.6$ $23.1$ $0.605$ $0.498$ $0.433$ $0.481$
ASR-Pretraining (52M) $26.5$ $12.2$ $1.82$ $9.00$ $51.9$ $32.8$ $16.4$ $27.1$ $0.680$ $0.537$ $0.443$ $0.511$
Hu-Transformer (113M) $24.3$ $11.4$ $2.18$ $8.60$ $49.9$ $31.9$ $17.0$ $26.8$ $0.650$ $0.522$ $0.439$ $0.499$
DSU-Adapter (48M) $26.5$ $12.9$ $1.76$ $9.13$ $52.1$ $33.9$ $16.5$ $27.4$ $0.681$ $0.548$ $0.442$ $0.513$
Table 1: Results in BLEU, chrF and COMET-22-DA on the test set of CoVoST-2 (X-En) by resource group. In all metrics, DSU-Adapter is much better than Hu-Transformer, which is $2.3$ times larger, in both "High" and "Mid" groups. DSU-Adapter, which does not requires transcripts in training, is also on a par with ASR-Pretraining. The best result in each group is denoted by ’_’.
5 Results and Analysis
Before discussing the results, it is worth noting that (1) Hu-Transformer is not memory-adaptive, and (2) ASR-Pretraining requires transcripts, unlike DSU which is self-supervised. Both methods are introduced for reference purposes of if such resources are available.
5.1 Improvement brought by DSU-Adapter
We divide the 21 language pairs by resource level into: 1) "High": {ca, de, es, fr}, 2) "Mid": {fa, it, pt, ru and zh}, 3) "Low": {ar, cy, et, id, ja, lv, mn, nl, sl, sv, ta, tr} and 4) "All": the 21 languages pairs. We report the average BLEU444nrefs:1—case:mixed—eff:no—tok:13a—smooth:exp—version:2.3.1 and chrF555nrefs:1—case:mixed—eff:yes—nc:6—nw:0—space:no—version:2.3.1 over the test sets of each group using sacrebleu Post (2018). In addition, we also provide the result in WMT22-COMET-DA Rei et al. (2022), which the source inputs are the gold-reference transcripts.
Table 1 compares our DSU-Adapter and the baselines. Our DSU-Adapter is $3$ BLEU (in the group "All") higher than the Scratch model. This shows that our proposed method of using DSU-pretraining can strengthen direct end-to-end ST without requiring transcripts and remain flexible in memory footprint (smaller in size than the HuBERT model). Furthermore, it is better than Hu-Transformer in spite of having half the parameters. For "Mid" and "High", the improvement in BLEU is $1.49$ and $2.23$ points respectively, but it falls short by $0.42$ points for "Low". We also compare to ASR pretraining, which is not always applicable, e.g., in low-resource setting or perhaps even in an unwritten language Zhang et al. (2022a). Surprisingly, our adapter is on a par with it, and its BLEU is $0.13$ points better. The result remains consistent when it is measured in chrF and COMET.
5.1.1 Language-specific performance
Figure 3 shows the performance on each language pair in BLEU, chrF and COMET-22-DA. Our DSU-Adapter (in green triangles) show consistent improvement over the Scratch model (in blue circles) in all language pairs. Such improvement is rather surprising since HuBERT-Base was trained solely on English audio data. We hypothesized that the cross-lingual improvement is related to HuBERT’s ability to capture language independent features, e.g. phonetic properties Pasad et al. (2023).
Compared with Hu-Transformer, DSU-Adapter maintains an evident improvement over most language pairs in both "High" and "Mid" groups. Exceptions are in "fa" and "pt", but the lags are almost negligible. In group "Low", Hu-Transformer is slightly better, especially in "nl" and "sv" pairs. However, most translation in this group is barely around $2$ BLEU, and the lags are small.
In most language pairs, DSU-Adapter performs similarly to ASR-pretraining (in red diamonds), except translating from "ru" audios. The improvement in this "ru-en" pair makes DSU-Adapter to have an evident advantage of $0.7$ BLEU in the group "Mid".
@@ 5.1.1 Language-specific performance
Table 2: (DSU) tokenization effect on 4 ST methods. Each ST model’s performance on the CoVoST-2 test set is measured by BLEU on group "All". All 4 methods could perform better than the Scratch model of $5.99$ BLEU as shown on Table 1. In general, darker (brighter) cells refer to weaker (stronger) models. The best two models apply both BPE on the DSU and separate vocabularies in PT (cells in yellow).
5.2 Tokenization effect to the DSU-to-Trl method and the DSU-Adapter method
In this section, we investigate how tokenization, including BPE, affects the DSU-to-Trl method and the DSU-Adapter method. We are particularly interested in their robustness toward the tokenization, especially using BPE on the DSU, since tuning the quantization process and retraining the subsequent models is computationally expensive.
In Table 2, the 1st column "Has BPE on DSU?" indicates if BPE is applied on the DSU. If "Yes", multiple DSU could be merged into one subword unit, e.g., ’#1 #456 #23 #999’ could be merged into ’#1#456#23#999’. The 2nd column "$|\mathcal{V}|$" shows the vocabulary configuration: its size, and if the model has a joint vocabulary. For example, "$1$K-$8$K" means that we use a vocabulary of size $1$K for DSU and a second vocabulary of size $8$K for English so that the DSU-to-Trl model would have separate vocabularies for the source (DSU) and target (English) sides. All results are in BLEU averaged over all language pairs, i.e., group "All".
5.2.1 DSU-to-Trl: robust to tokenization?